Home‎ > ‎Reports‎ > ‎Crew100B‎ > ‎

day07


Crew 100B Log Book for March 4 2011

Photos of the day





Commander's Check-In Report

Balwant Rai reporting

Experimental Status: All the medical, habitability, astrobiological and food experiments began today with unique results.

Time Departed/Returned from EVA: three EVAs performed

Time Departed/Returned from EVA: As per EVA reports

The meals today were just satisfying since it was a non cooking day.

EVA’s: We completed three successful EVA’s.

Number Of reports: 3 science reports, 1 engineering report, 1 journalist report, 1 geologist report, 3 EVA reports and 1 chef report.

Plan for tomorrow: will begin the massages to lower the surface tensions and then medical experiment followed by habitability again on all crew members. Interpretation of the Sun by means of the radio telescope will also be done. The study of the rover’s maneuverability as well as the potential use of the drone, weather pending, will take place. Soil samples will yet again be collected during three EVA’s tomorrow. One special EVA will be planned to learn the emergency objectives as planned by myself in rotation here in MDRS crew 78 before in 2009. One will be on foot, and two will use the ATV’s. Tomorrow will be a cooking day.


Commander's Report

Balwant Rai Reporting

Today we commenced the day with our previous crew experiments on habitability and then continuing medical experiments and our habitability experiment.

The crew’s medical health will continue to be monitored by our health and safety officer for the remainder of our stay here on Mars. Breakfast was satisfying. Our crew had their saliva samples taken (via Versi Sal 1 from Oasis Diagnostics), vital parameters monitored (via the Zephyr Bioharness), and their heart rate variability measured (via NERV Express 4.2 software). I also performed the cognitive study (via CogState Research software) on all crew members, as well. We began with our new experiment of sensory test and affects on CNS that comprised of natural fragrances and had no artificial one to make sure crew members feel homely and contended( also conducted in crew 78 with good results in 2009).

Our crew scientist continued to work through problems with starting observations of the sun via the radio telescope. Three EVA’s were performed to sites of interest. Today is a non-cooking day, so our chef is planning to make something filling for all crew members. Tomorrow, we will continue our experiments starting massages suggested by our HSO for all crew members to have proper sleep as some of us experienced little problems. Medical, habitability and astrobiological experiments will be taken into account, also, the rover experiment and the radio telescope experiments will be performed.


Engineering Report

Matthieu Ansart Reporting

Generator/Electricity:
· Kitty (Diesel Generator) (Used? Yes/No): Yes
· Honey (Gas Generator) (Used? Yes/No): No
· Battery Charge Level: Checked by DG
o 10.59 V everstart
o 12.37 V armor plate
o 13.32 V green
Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good

Fuel Status:
· Diesel (Full/0.75/0.5/0.25/Empty): between 0.75 and 0.5
· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.3
· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 1,5 containers full
Notes/Comments:



Oil:
· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG
· Honey Oil Quantity (# of Quarts in storage): Leave it to DG
· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench
· Car Oil Quantity (# of Quarts in storage):



Water Status:


· Outside Potable Water Tank Level (inches from bottom): 10


· Trailer Potable Water Tank Level (inches from bottom): 1/3 Full


· Hab Potable Water Tank Level (inches from bottom): 12


· Potable Water Meter Reading:



Notes/Comments : No leak after the using of the heat water today.



GreenHab:


· Condition of plants in Tank 1: green water


· Condition of plants in Tank 2: green plants and snails


· Condition of plants in Tank 3: green plants


· Greenhab notes/ comments:




Observatory:


· Telescope Used? (Yes/no) No



Observatory Notes:


Transportation:



Hab Car (NOT RENTAL)


· Used (yes/no): No


· Oil Checked (yes/no): No


· Tire Status: Not checked


· Odometer at end of day: 123973


· Notes/Comments on Hab Car:



ATV 1 (Opportunity):


· Used: Yes


· Oil Checked: no


· Fuel Consumed:Replaced


· Tire Status: good


· Comments:



ATV 2 (Spirit):


· Used: No


· Oil Checked: No


· Fuel Consumed:


· Tire Status:


· Comments: It has been taken back.



ATV 3 (Viking I):
· Used: Yes
· Oil Checked: no
· Fuel Consumed: Replaced
· Tire Status: good
· Comments:
ATV 4 (Viking II):
· Used: Yes
· Oil Checked: no
· Fuel Consumed:Replaced
· Tire Status: good
· Comments:

Heating and Ventilation:
· Thermostat setting upstairs (Farenheit): 68
· Actual temperature upstairs (Farenheit): 68
· Thermostat setting downstairs (Farenheit): 55
· Actual temperature downstairs (Farenheit): 61
Computers/Networking Infrastructure:
Notes/comments on internet/computers?:
General Engineering

1. Notes/Comments:
Clear


Science Report (Medical)

Jasdeep Kaur Reporting

Time: 0930-Blood pressure and pulse rate monitored by Zephyr technology and by manually prior and after experiments.

Crew Members: Science Team: 6 crew members Balwant Rai, Rachel Dompnier , Quentin Bourges,Chrystal Latham,Matthieu Ansart and Jasdeep Kaur

Location: Hab

Objective: To examine the affects of different activities on blood pressure, heart rate and pulse rate.

Observation: All crew members were monitored for their blood pressures. The blood pressures and heart rate was also monitored by Zephyr technology. Pulse was checked manually and electronically. This includes all IVA and EVA activities

Results:There is a difference in blood pressures examined after EVA activity in all crew members, Also, heart rate varies with EVA and IVA activities but in safe limits.

Lessons learned: Some of other massages to relieve surface tensions will be taught in coming days to be motivated and balance the body functions.


Science Report (Astronomy)

Crystal Latham Reporting

The main goal of my research here at MDRS is to monitor the Sun for solar storms using the decametric Radio telescope. A few days ago I was able to get the radio receiver working with a simple fix, using tape to hold the volume knob in place, so that a signal could be heard. However, today the receiver decided this quick fix was not acceptable.

After discussing the problem with a number of people who have experience with radio telescopes, like the one here at MDRS, a possible source for the problem has been located. I did a number of measurements of voltage across capacitors and jumpers that are related to the volume knob/power switch, and found that the jumper did not have voltage going across it when the power was turned on. The capacitor, closest to the volume knob, did have the correct about of voltage going across it, both when the receiver was off and on.

This is a work and progress, and I can only hope that a solution will be found soon to fix the radio receiver. In the meantime, since I have received the visual astronomer box, I will have an observing session tonight using the Musk Observatory. This session will include getting the telescope aligned properly and observing the Orion Nebula, the Pleiades, and perhaps some other deep-sky objects. Some of the crew is very interested in seeing the Musk observatory in action, since some have never even looked through a telescope before.


Science Report (Geologist)

Jasdeep Kaur Reporting

Time, 14:10-15:00

Crew Members: Crystal Lantham, Jasdeep Kaur, Balwant Rai

Site Location: Confidential

Transit Mode: ATVs

Objectives: Geo-Mission: to collect samples by following all protocols

Results: This study is especially performed to study the contamination, so measures can be taken to prevent the planet MARS. Samples were collected to for contamination study in a particular area with characteristics of water flow in past. Four samples were taken from same area with same distance from each other. First the test-tubes were placed equidistant from each other and then samples were collected in the same test-tubes following all protocols.

Lessons Learned: To be careful in placing test-tubes equi-distant.


Science Report (Human Factors)

Balwant Rai and Jasdeep Kaur Reporting
JBR study of the Human Dimension and factors: MDRS Crew 100 B ILEWG Euro MoonMars crew
In long duration space mission, these two bi-directional targets are discrete for the technical system and the human system in a unique way. The human aspects are imperative for a human space mission to achieve the successful operation. As the technical system is less complex as compared to the human system, the focus is ought to be on human system. Habitability simulations engaged in MDRS Crew 100 B ILEWG Euro Moon Mars crew determined from standardized JBR questionnaire from all crew members whether the design of the physical living environment fulfils primary necessities of the human being and has the competence within a specific space analogous environment as on MARS to keep up a certain quality of life. This take into account the simulation of requirements for the usability of space (living, working and movement), the support requirements (medical, food, and life support) as well as characterizing fundamental quantities and qualities definitions of the requirements. Eating jointly seems to be noteworthy for the crew and also they yet place up common meals to socialize in this way and to keep the collective tensions little. I will put forward this routine to bring ahead positive results of missions. Listening to music and watching movies together was a tool to build up motivated team. According to survey, the division of living and working area would give more comfortable results. The crew quarters were accepted by the crewmembers with minor modifications. One of our crew member had problem in sleeping because of cold environment. It might be due to lack of proper sleeping bag, also could be due to certain air passage connections.

EVA 19 Report

Quentin Bourges Reporting

EVA#19 09:30 am to 10:30 am

Crew members: Rachel Dompnier, Matthieu Ansart, Quentin Bourges

Location : Green hab and near by

Transit mode: Foot

Objective: Try the new Virtual Reality glasses with the UAV.

Results: We had to make some settings and only one Ipad software on three could work with the VR glasses. These glasses allowed us to see the UAV’s front video camera as we were piloting it. This can fix the problem we had with the reflection of the sun on the iPad that impeded us to see where we were going with the UAV.

Lessons learned: When we have the glasses on, we can see the video camera of the UAV in our field of view. This is very tricky because the angle of vision is much reduced even if we don’t have this impression. So it is easy to knock against a corner. That’s why we have to keep practicing. 


EVA 20 Report

Jasdeep Kaur Reporting

EVA # 20, 14:10-15:00

Crew Members: Crystal Lantham, Jasdeep Kaur, Balwant Rai

Site Location: Confidential

Transit Mode: ATVs

Objectives: Geo-Mission: to collect samples by following all protocols

Results: Samples were collected for contamination study. Four samples were taken from same area at same distance from each other. First the test-tubes were placed equidistant from each other and then samples were collected in the same test-tubes following all protocols.

Lessons Learned: To always use gloves while driving the ATVs


EVA 21 Report

Rachel Dompnier Reporting
EVA #21, 3:10 to 4:40 pm
Crew Members: Matthieu Ansart, Quentin Bourgès, Rachel Dompnier
Site Location: N 38.38960° / W110.82346°
Transit Mode: ATV’s
Objectives: Take soil samples. Record active blood pressure.
Results: We found some stones with crystal properties but we do not know what it exactly is.


Chef's Report

Rachel Dompnier Reporting

1a) Today was a non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

No

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Santa Fe Black Beans with rice and chicken

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

No leftover.

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

Nothing

5) Lessons learned (if any)

I was scared of trying to put water on the dry food instead of putting the food in the water but I think it is actually better. The food does not stick and it is well cooked.

6) Comments/questions for Kim and Jean, the food study investigators

Nothing


Journalist Report

Crystal Latham Reporting

Today was crew 100B’s seventh day on Mars. In the past week crew 100 B has accomplished a lot, including having a total of 21 EVA’s! In the past week, crew members have learned to drive a rover, collect samples (the proper way), and use complicated electronic equipment. With eight medical experiments a day to be done, there is never a moment of down time.

Today there were three very successful EVA’s, which resulted in interesting data for both our drone, medical, and astrobiological experiments. Interesting rock samples, were collected, that look like petrified wood. There were also some samples of what looks to be crystal or quartz. Since no one on crew 100 B is a certified geologist, we will have to wait for experts to confirm these discoveries. Soil samples were also taken, in areas that indicated water had been flowing.

The amount crew 100 B is doing is starting to weigh down some of the crew. Many miss their families and friends, and can only hope for a spare moment to send an email to say hi. Sleep is something that many members of the crew are missing because of many factors. Every night one hopes that tonight will be the night for a decent sleep.

Tomorrow marks the weekend here on Mars, but no parties or breaks will be had. It’s all work, work, work here on Mars, and maybe two minutes for your own thoughts prior to sleep. Crew 100 B is wishing all those Earthlings reading this a fun and joyful weekend.